OpenStudy (anonymous):

Write as a single natural logarithm: 12ln(x)-ln(y)-ln(z)

OpenStudy (anonymous):

ln(a) - ln(b) = ln(a/b)

OpenStudy (anonymous):

\(ln(a^x) = x*ln(a)\)